OpenStudy (jazzyfa30):

Simplify: (6y2 -6/ 8y2 +8y)/(3y -3 / 4y2 + 4y)

OpenStudy (campbell_st):

well the rule of thumb for dividing fractions is flip the denominator and multiply so you get \[\frac{6y^2 - 6y}{8y^2 + 8y} \times \frac{4y^2 + 4y}{3y^2 - 3y}\] a little factoring will enable you to cancel common factors and find the solution.

OpenStudy (jazzyfa30):

?????? can u walk me through it

OpenStudy (jazzyfa30):

@phi

OpenStudy (campbell_st):

well here is a little simplification \[\frac{2(3y^2 - 3y)}{2(4y^2 + 4)} \times \frac{4y^2 +4}{3y^2 -3y}\] what every you see in the numerator and is also in the denominator can be cancelled.

OpenStudy (jazzyfa30):

????????

OpenStudy (campbell_st):